# On Edge Coloring of Multigraphs

### v4

@@ 3 Extented Tashkinov Tree (ETT)
(A2) Let $n$ be a nonnegative integer. Suppose (A1) holds for all ETTs with at most $n$ rungs, and (A2), (A2.5), and (A3) hold for all ETTs with at most $(n-1)$ rungs. If $T_{n+1}$ is a closed ETT with ladder $T_{0}\subset\dots\subset T_{n}\subset T_{n+1}$, then $T_{n+1}$ has the interchangeability property. Furthermore, if $T_{n+1}$ is $RE$ finished, then any $(T_{n+1},D_{n},\varphi)$-stable coloring $\varphi^{\prime}$ is a $T_{n+1}$ mod $\varphi$ coloring, and $T_{n+1}$ remains $RE$ finished under $\varphi^{\prime}$.
@@
Let $T_{n+1}$ be a closed ETT with ladder $T_{0}\subset\dots\subset T_{n}\subset T_{n+1}$ and $h$ be the initial index of $T_{n+1}$. We first prove the “Furthermore” part. Let $\varphi^{\prime}$ be a $(T_{n+1},D_{n},\varphi)$-stable coloring and assume that $T_{n+1}$ is $RE$ finished. If $n=0$, then $T_{1}$ remains a closed Tashkinov tree under $\varphi^{\prime}$ and we have as desired. So we assume $1\leq h\leq n$. Since $T_{h}$ is elementary under $\varphi$ by (A1), there is exactly one vertex $v\in V(T_{h})$ with $\gamma_{h}\in\overline{\varphi}(v)$ and $\delta_{h}\notin\overline{\varphi}(T_{h})$ as $T_{h}$ is closed under $\varphi$ (see Algorithm 3.1). Therefore, by (A2) on $T_{h}$ which is $RE$ finished and has $h-1<n$ rungs, there is exactly one $(\gamma_{h},\delta_{h})$-path intersecting $T_{h}$ under $\varphi$. Let $u$ and $f$ be the corresponding $(\gamma_{h},\delta_{h})$-exit vertex and $(\gamma_{h},\delta_{h})$-exit edge, respectively. Then under $\varphi$, all the $\delta_{h}$ colored boundary edges of $T_{h}$ are contained some $(\gamma_{h},\delta_{h})$-ear of $T_{h}$, except for $f$. Since $T_{n+1}$ is $RE$ finished, all the $(\gamma_{h},\delta_{h})$-ears of $T_{h}$ are contained in $G[V(T_{n+1})]$ (see Algorithm 3.1), and therefore they are colored the same under $\varphi^{\prime}$ as $\varphi$, because $\varphi^{\prime}$ is $(T_{n+1},D_{n},\varphi)$-stable and $\{\gamma_{h},\delta_{h}\}=S_{h}\subseteq\overline{\varphi}(T_{h})\cup D_{h}\subseteq\overline{\varphi}(T_{n+1})\cup D_{n}$. This forces $u$ and $f$ be the only $(\gamma_{h},\delta_{h})$-exit vertex and $(\gamma_{h},\delta_{h})$-exit edge of $T_{h}$ under $\varphi^{\prime}$, respectively, and all the $(\gamma_{h},\delta_{h})$-ears of $T_{h}$ under $\varphi^{\prime}$ are contained in $G[V(T_{n+1})]$. So by (A2.5) with $T_{n}$ that has $(n-1)$-rungs in place for $T_{n+1}$, the $(T_{n+1},D_{n},\varphi)$-stable coloring $\varphi^{\prime}$ is also $T_{n}+f_{n}$ mod $\varphi$. Thus any tree-sequence obtained from $T_{n}+f_{n}$ is an ETT under $\varphi^{\prime}$ with $n$ rungs. Since $\varphi^{\prime}$ is $(T_{n+1},D_{n},\varphi)$-stable and $T_{n+1}$ is closed under $\varphi$, we see that $T_{n+1}$ remains a closure of $T_{n}+f_{n}$ under $\varphi^{\prime}$, so $\varphi^{\prime}$ is $T_{n+1}$ mod $\varphi$. Moreover, since all the $(\gamma_{h},\delta_{h})$-ears of $T_{h}$ under $\varphi^{\prime}$ are contained in $G[V(T_{n+1})]$, $T_{n}$ is also $RE$ finished under $\varphi^{\prime}$. We now prove that $T_{n+1}$ has the interchangeability property. Let $\alpha,\beta$ be two colors such that $\alpha\in\overline{\varphi}(T_{n+1})$, and furthermore $\alpha\neq\gamma_{n}$ if $T_{n+1}$ is not $RE$ finished. We first show that $T_{n+1}$ has the interchangeability property by assuming otherwise that there are two $(\alpha,\beta)$-paths $Q_{1}$ and $Q_{2}$ intersecting $T_{n+1}$. By (A1), $T_{n+1}$ is elementary with respect to $\varphi$. So $T_{n+1}$ contains at most two vertices $v$ with $\overline{\varphi}(v)\cap\{\alpha,\beta\}\neq\emptyset$, which in turn implies that at least two ends of $Q_{1}$ and $Q_{2}$ are outside $T_{n+1}$. Since $T_{n+1}$ is closed, $\beta\notin\overline{\varphi}(T_{n+1})$. Thus we further deduce that at least three ends of $Q_{1}$ and $Q_{2}$ are outside $T_{n+1}$. Traversing $Q_{1}$ and $Q_{2}$ from these ends respectively, we can find at least three $(\alpha,\beta)$-exit paths $P_{1},P_{2},P_{3}$ for $T_{n+1}$ under $\varphi$ with corresponding exit vertices $b_{1},b_{2}$ and $b_{3}$. For $i=1,2,3$, let $a_{i}$ be the end of $P_{i}$ outside of $T_{n+1}$, and $f_{i}$ be the boundary edge of $T_{n+1}$ in $P_{i}$. Renaming subscripts if necessary, we may assume that $b_{1}\prec b_{2}\prec b_{3}$ along $T_{n+1}$. Define $Q(T_{n+1})=0$ if $\delta_{h}\in\overline{\varphi}(T_{n+1}(b_{2}))$, and $Q(T_{n+1})=1$ otherwise. So $Q(T_{1})=0$ as $D_{0}=\emptyset$ for any ETT $T_{1}$ with 00-rungs. We call the tuple $(\varphi,T_{n+1},\alpha,\beta,P_{1},P_{2},P_{3})$ a counterexample and use ${\cal K}$ to denote the set of all such counterexamples. With a slight abuse of notation, let $(\varphi,T_{n+1},\alpha,\beta,P_{1},P_{2},P_{3})$ be a counterexample in ${\cal K}$ such that:
@@
Indeed, if $T_{n+1}$ is $RE$ finished, then we have as claimed by the “furthermore” part we proved earlier. Now if $T_{n+1}$ is not $RE$ finished, we have $\alpha\neq\gamma_{h}$ by assumption and $\delta_{n}\notin\overline{\varphi}(T_{n+1})$. So both $\gamma$ and $\alpha$ are different from $\gamma_{h}=\gamma_{n}$ and $\delta_{n}$, and therefore $T_{h}$ has the same $(\gamma_{h},\delta_{h})$-exit vertex under $\varphi_{\gamma}$ as $\varphi$. So $\varphi_{\gamma}$ is $T_{n}+f_{n}$ mod $\varphi$ by (A2.5) because $\varphi_{\gamma}$ is also $(T_{n},D_{n},\varphi)$-stable. Thus any closure of $T_{n}+f_{n}$ is an ETT with $n$-rungs under $\varphi_{\gamma}$. Finally since $\varphi_{\gamma}$ is $(T_{n+1},D_{n},\varphi)$-stable, we have as claimed as $T_{n+1}$ remains a closure of $T_{n}+f_{n}$ under $\varphi_{\gamma}$. Furthermore, we have
@@
Note that (2) holds trivially if $n=0$, so we may assume $n\geq 1$, and therefore $h\geq 1$. Since $|\overline{\varphi}(T_{1})|\geq 5$ by Lemma 3.2, there’s a color $\gamma\in\overline{\varphi}(T_{n})$ with $\gamma\neq\gamma_{h}$. So by (1), $\varphi_{\gamma}=\varphi/(G-T_{n+1},\alpha,\gamma)$ is $T_{n+1}$ mod $\varphi$ and $T_{n+1}$ is an ETT with $n$-rungs under $\varphi_{\gamma}$. Thus $T_{n}$ remains an ETT with $(n-1)$-rungs under $\varphi_{\gamma}$. Since $\gamma\neq\gamma_{h}$, we see that there is at most one $(\gamma,\beta)$-path intersecting $T_{n}$ by (A2). Because $P_{1}$, $P_{2}$, and $P_{3}$ are $(\gamma,\beta)$-exit paths of $T_{n+1}$ under $\varphi_{\gamma}$ with exit vertices $b_{1}\prec b_{2}\prec b_{3}$, (2) is established. Depending on whether $\beta$ is $\delta_{h}$ or not, we consider the following two cases.
Case I:
$\beta\neq\delta_{h}$.
@@
Note that $P_{3}$ is a $(\gamma,\beta)$-path under $\varphi_{\gamma}$ with $\gamma\in\overline{\varphi}_{\gamma}(b_{3})$. So $f_{i}$ is colored by $\beta$ under both $\varphi$ and $\varphi_{\gamma}$ for $i=1,2,3$. Consider $\sigma=\varphi_{\gamma}/P_{3}$. Note that $\beta\in\overline{\sigma}(b_{3})$, and $f_{i}$ is colored by $\beta$ under $\sigma$ for $i=1,2$. Let $T^{\prime}$ be a closure of $T_{n+1}(b_{3})$ under $\sigma$. Since $\beta\in\overline{\sigma}(b_{3})$, by TAA, (4) both $f_{1}$ and $f_{2}$ are contained in $G[V(T^{\prime})]$. We claim that (5) $\sigma$ is $T_{n}+f_{n}$ mod $\varphi_{\gamma}$, and $T^{\prime}$ is an ETT with $(n-1)$-rungs under $\sigma$. Note that $T^{\prime}$ can be obtained by TAA from $T_{n}+f_{n}$. So by (1) and (3), $T^{\prime}$ is an ETT with $(n-1)$-rungs under $\sigma$ once we show that $\sigma$ is $T_{n}+f_{n}$ mod $\varphi_{\gamma}$. Indeed, because $\beta\notin\overline{\varphi}(T_{n+1})$ and $T_{n+1}$ is closed under $\varphi$, we have $\beta\neq\gamma_{h}$. As $\gamma_{h}\in\overline{\varphi}(T_{h})$ and $T_{n+1}$ is elementary under $\varphi$ by (A1), we have $\gamma\neq\gamma_{h}$ by (2). So both $\gamma$ and $\beta$ are different from $\gamma_{h}$ Recall that we have $\beta\neq\delta_{h}$ in the current case. So if $\gamma\neq\delta_{h}$, then $T_{h}$ has the same $(\gamma_{h},\delta_{h})$-exit vertex under $\sigma$ as $\varphi_{\gamma}$, and thus $\sigma$ is $T_{n}+f_{n}$ mod $\varphi_{\gamma}$ by (A2.5) for $T_{n}$. If $\gamma=\delta_{h}$, then all the $(\gamma_{h},\delta_{h})$-ears of $T_{h}$ under $\varphi$ are contained in $G[V(T_{n+1})]$ by TAA. Thus by the construction of $\varphi_{\gamma}$ and $\sigma$, they are colored the same under $\sigma$ as under $\varphi$. Since $T_{h}$ only has one $(\gamma_{h},\delta_{h})$-exit path under $\varphi$ by (A2), $T_{h}$ still has the same $(\gamma_{h},\delta_{h})$-exit vertex under $\varphi_{\gamma}$ as $\varphi_{\gamma}$, and thus (5) is again established by (A2.5) for $T_{n}$.
@@
Case II:
$\beta=\delta_{h}$.
@@
Observe that $P_{j}$ is a $(\eta,\beta)$-path under $\varphi_{\eta}$ with $\eta\in\overline{\varphi}_{\eta}(b_{j})$. Similarly as in the previous case, we let $\sigma=\varphi_{\eta}/P_{j}$ and $T^{\prime}$ be a closure of $T_{n+1}(b_{j})$ under $\sigma$. We claim that
@@
Note that following (7), $T^{\prime}$ is elementary by (A1) and (6), and therefore similarly as earlier, none of $a_{1}$ and $a_{j}$ are in $T^{\prime}$ (see the paragraph before Case II). So under $\sigma$, $P_{1}$ and $P_{j}$ guarantee two $(\beta,\eta)$-paths intersecting $T^{\prime}$, and they provide three $(\beta,\eta)$-exit paths $L_{1},L_{2}$ and $L_{3}$ of $T^{\prime}$, where $L_{1}$, $L_{2}$, and $L_{3}$ produce at least two $(\beta,\eta)$-exit vertices in $T^{\prime}$ not before $b_{j}$ along $\prec$ of $T^{\prime}$ (one of them might be $b_{j}$ itself). Since $\delta_{h}=\beta\in\overline{\sigma}(b_{j})$, $(\sigma,T^{\prime},\beta,\eta,L_{1},L_{2},L_{3})$ provides a counter-example with $Q(T^{\prime})=0$, a contradiction.
@@
We now prove (7) to finish the proof of Lemma 4.1 by assuming otherwise that $T_{h}$ has a different $(\gamma_{h},\delta_{h})$-exit vertex $v_{1}$ under $\sigma$ from the $(\gamma_{h},\delta_{h})$-exit vertex $v_{2}$ under $\varphi_{\eta}$. Note that $\sigma$ is $(T_{n},D_{n},\varphi_{\eta})$-stable as $P_{j}$ is disjoint from $T_{n}$ (see (2)). So $\sigma$ is $T_{h}$ mod $\varphi$ and $T_{h}$ has exactly one $(\gamma_{h},\delta_{h})$-exit vertex $v_{1}$ under $\sigma$ by (A2) for $T_{h}$ with $h\leq n$. We first assume that $v_{1}\prec v_{2}$ along $\prec$ of $T_{n+1}$. In this case, under $\varphi_{\eta}$, $P_{j}$ intersects both the $(\gamma_{h},\delta_{h})$-exit path $P_{ex}$ and a $(\gamma_{h},\delta_{h})$-ear $D$ with root $v_{1}$ of $T_{h}$. Since $\gamma\neq\gamma_{h}$ and $\eta\neq\gamma_{h}$ (see the sentence before (6)), $P_{ex}$ is still a $(\gamma_{h},\delta_{h})$-exit path and $D$ is still a $(\gamma_{h},\delta_{h})$-ear of $T_{h}$ under $\varphi_{\gamma}$. Let $n^{\prime}$ be the smallest index such that $v_{2}\in T_{n^{\prime}}$ and $T_{n^{\prime}}$ is $RE$ finished under $\varphi_{\gamma}$. Then $n^{\prime}\leq n$ and $v_{2}\notin T_{h^{\prime}}$, where $h^{\prime}$ is the initial index of $n^{\prime}$. Moreover, under $\varphi_{\gamma}$, $P_{ex}$ is a $(\gamma_{h},\delta_{h})$-exit path of $T_{n^{\prime}}$ by our choice of $n^{\prime}$. Since $T_{n^{\prime}}$ is an ETT with at most $(n-1)$-rungs by (1), $v_{1}$ is a root of a $(\gamma_{h},\delta_{h})$-ear $D^{\prime}$ of $T_{n^{\prime}}$ by (A2). Since $v_{1}$ is a root of the $(\gamma_{h},\delta_{h})$-ear $D$ of $T_{h}$ with $h\geq n^{\prime}$, $D$ is a subpath of $D^{\prime}$. So $P$ is a $(\gamma,\delta_{h})$-path intersecting both $P_{ex}$ and $D^{\prime}$. Note that $\gamma\in\overline{\varphi}_{\gamma}(T_{n^{\prime}})$ as $x_{e}$ is the first vertex in $T_{n+1}$ along $\prec$. However, this is a contradiction to the later introduced technical Lemma 4.4 under $\varphi_{\gamma}$ with $T_{n^{\prime}}$ ($n^{\prime}\leq n$) in place of $T_{n+1}$, $D^{\prime}$ in place of $D$, $v_{2}$ in place of $x$, $v_{1}$ in place of $y$, $\gamma$ in place of $\alpha$, and $\gamma_{h},\delta_{h}$ in place of $\gamma,\delta$.
@@
So we may assume $v_{2}\prec v_{1}$. Now in this case, under $\sigma$, with slightly abuse of notations, we may assume $P_{j}$ intersects both the $(\gamma_{h},\delta_{h})$-exit path $P_{ex}$ of $T_{h}$ with exit $v_{1}$ and a $(\gamma_{h},\delta_{h})$-ear $D$ with root $v_{2}$. Since $T_{n+1}$ is elementary under $\varphi$, $\delta_{h}=\beta\notin\overline{\varphi}(T_{n+1})$, and $\gamma,\eta\neq\gamma_{h}$ (see the sentence before (6)), we observe that:
@@
(8) $\sigma$ is $(T_{h},D_{h},\varphi)$-stable, $T_{n+1}$ is elementary under $\sigma$, $\sigma(P)=\{\gamma,\eta,\delta_{h}\}$, all the $\eta$-colored edges of $P$ are not in $G[V(T_{n+1})]$, all the $\gamma$ colored edges of $P$ are in $G[V(T_{n+1})]$, $\delta_{h}\in\overline{\sigma}(b_{j})$, two ends of $P$ miss either $\delta_{h}$ or $\eta$, and exactly one boundary edge of $T_{n+1}$ is colored by $\eta$ which is $f_{j}$ under $\sigma$.

### v5

@@ 3 Extented Tashkinov Tree (ETT)
(A2) Let $n$ be a nonnegative integer. Suppose (A1) holds for all ETTs with at most $n$ rungs, and (A2), (A2.5), and (A3) hold for all ETTs with at most (n1)(n-1) rungs. If $T_{n+1}$ is a closed ETT with ladder $T_{0}\subset\dots\subset T_{n}\subset T_{n+1}$, then $T_{n+1}$ has the interchangeability property. Furthermore, if $T_{n+1}$ is $\operatorname{RE}$ finished, then any $(T_{n+1},D_{n},\varphi)$-stable coloring $\varphi^{\prime}$ is a $T_{n+1}$ mod $\varphi$ coloring, and $T_{n+1}$ remains $\operatorname{RE}$ finished under $\varphi^{\prime}$.
@@
Let $T_{n+1}$ be a closed ETT with ladder $T_{0}\subset\dots\subset T_{n}\subset T_{n+1}$ and $h$ be the initial index of $T_{n+1}$. We first prove the “Furthermore” part. Let $\varphi^{\prime}$ be a $(T_{n+1},D_{n},\varphi)$-stable coloring and assume that $T_{n+1}$ is $\operatorname{RE}$ finished. If $n=0$, then $T_{1}$ remains a closed Tashkinov tree under $\varphi^{\prime}$ and we have as desired. So we assume $1\leq h\leq n$. Since $T_{h}$ is elementary under $\varphi$ by (A1), there is exactly one vertex $v\in V(T_{h})$ with $\gamma_{h}\in\overline{\varphi}(v)$ and $\delta_{h}\notin\overline{\varphi}(T_{h})$ as $T_{h}$ is closed under $\varphi$ (see Algorithm 3.1). Therefore, by (A2) on $T_{h}$ which is $\operatorname{RE}$ finished and has h1<nh-1<n rungs, there is exactly one $(\gamma_{h},\delta_{h})$-path intersecting $T_{h}$ under $\varphi$. Let $u$ and $f$ be the corresponding $(\gamma_{h},\delta_{h})$-exit vertex and $(\gamma_{h},\delta_{h})$-exit edge, respectively. Then under $\varphi$, all the $\delta_{h}$ colored boundary edges of $T_{h}$ are contained in some $(\gamma_{h},\delta_{h})$-ear of $T_{h}$, except for $f$. Since $T_{n+1}$ is $\operatorname{RE}$ finished, all the $(\gamma_{h},\delta_{h})$-ears of $T_{h}$ are contained in $G[V(T_{n+1})]$ (see Algorithm 3.1), and therefore they are colored the same under $\varphi^{\prime}$ as $\varphi$, because $\varphi^{\prime}$ is $(T_{n+1},D_{n},\varphi)$-stable and $\{\gamma_{h},\delta_{h}\}=S_{h}\subseteq\overline{\varphi}(T_{h})\cup D_{h}\subseteq\overline{\varphi}(T_{n+1})\cup D_{n}$. This forces $u$ and $f$ be the only $(\gamma_{h},\delta_{h})$-exit vertex and $(\gamma_{h},\delta_{h})$-exit edge of $T_{h}$ under $\varphi^{\prime}$, respectively, and all the $(\gamma_{h},\delta_{h})$-ears of $T_{h}$ under $\varphi^{\prime}$ are contained in $G[V(T_{n+1})]$. So by (A2.5) with $T_{n}$ that has (n1)(n-1)-rungs in place for $T_{n+1}$, the $(T_{n+1},D_{n},\varphi)$-stable coloring $\varphi^{\prime}$ is also $T_{n}+f_{n}$ mod $\varphi$. Thus any tree-sequence obtained from $T_{n}+f_{n}$ is an ETT under $\varphi^{\prime}$ with $n$ rungs. Since $\varphi^{\prime}$ is $(T_{n+1},D_{n},\varphi)$-stable and $T_{n+1}$ is closed under $\varphi$, we see that $T_{n+1}$ remains a closure of $T_{n}+f_{n}$ under $\varphi^{\prime}$, so $\varphi^{\prime}$ is $T_{n+1}$ mod $\varphi$. Moreover, since all the $(\gamma_{h},\delta_{h})$-ears of $T_{h}$ under $\varphi^{\prime}$ are contained in $G[V(T_{n+1})]$, $T_{n}$ is also $\operatorname{RE}$ finished under $\varphi^{\prime}$.
Let $\alpha,\beta$ be two colors such that $\alpha\in\overline{\varphi}(T_{n+1})$, and furthermore $\alpha\neq\gamma_{n}$ if $T_{n+1}$ is not $\operatorname{RE}$ finished. We now show that $T_{n+1}$ has interchangeability property by assuming otherwise that there are two $(\alpha,\beta)$-paths $Q_{1}$ and $Q_{2}$ intersecting $T_{n+1}$. By (A1), $T_{n+1}$ is elementary with respect to $\varphi$. So $T_{n+1}$ contains at most two vertices $v$ with $\overline{\varphi}(v)\cap\{\alpha,\beta\}\neq\emptyset$, which in turn implies that at least two ends of $Q_{1}$ and $Q_{2}$ are outside $T_{n+1}$. Since $T_{n+1}$ is closed, $\beta\notin\overline{\varphi}(T_{n+1})$. Thus we further deduce that at least three ends of $Q_{1}$ and $Q_{2}$ are outside $T_{n+1}$. Traversing $Q_{1}$ and $Q_{2}$ from these ends respectively, we can find at least three $(\alpha,\beta)$-exit paths $P_{1},P_{2},P_{3}$ for $T_{n+1}$ under $\varphi$ with corresponding exit vertices $b_{1},b_{2}$ and $b_{3}$. For $i=1,2,3$, let $a_{i}$ be the end of $P_{i}$ outside of $T_{n+1}$, and $f_{i}$ be the boundary edge of $T_{n+1}$ in $P_{i}$. Renaming subscripts if necessary, we may assume that $b_{1}\prec b_{2}\prec b_{3}$ along $T_{n+1}$. Define $Q(T_{n+1})=0$ if $\delta_{h}\in\overline{\varphi}(T_{n+1}(b_{2}))$, and $Q(T_{n+1})=1$ otherwise. So $Q(T_{1})=0$ as $D_{0}=\emptyset$ for any ETT $T_{1}$ with 0-rungs. We call the tuple $(\varphi,T_{n+1},\alpha,\beta,P_{1},P_{2},P_{3})$ a counterexample and use $\cal K$} to denote the set of all such counterexamples.
Let $(\varphi,T_{n+1},\alpha,\beta,P_{1},P_{2},P_{3})$ be a counterexample in $\cal K$} such that:
@@
Indeed, if $T_{n+1}$ is $\operatorname{RE}$ finished, then we have as claimed by the “Furthermore” part we proved earlier. Now if $T_{n+1}$ is not $\operatorname{RE}$ finished, we have $\alpha\neq\gamma_{h}$ by assumption and $\delta_{n}\notin\overline{\varphi}(T_{n+1})$. So both $\gamma$ and $\alpha$ are different from $\gamma_{h}=\gamma_{n}$ and $\delta_{n}$, and therefore $T_{h}$ has the same $(\gamma_{h},\delta_{h})$-exit vertex under $\varphi_{\gamma}$ as $\varphi$. So $\varphi_{\gamma}$ is $T_{n}+f_{n}$ mod $\varphi$ by (A2.5) because $\varphi_{\gamma}$ is also $(T_{n},D_{n},\varphi)$-stable. Thus any closure of $T_{n}+f_{n}$ is an ETT with $n$-rungs under $\varphi_{\gamma}$. Finally since $\varphi_{\gamma}$ is $(T_{n+1},D_{n},\varphi)$-stable, we have as claimed as $T_{n+1}$ remains a closure of $T_{n}+f_{n}$ under $\varphi_{\gamma}$. Furthermore, we have
@@
Note that (2) holds trivially if $n=0$, so we may assume $n\geq 1$, and therefore $h\geq 1$. Since $|\overline{\varphi}(T_{1})|\geq 5$ by Lemma 3.2, there’s a color $\gamma\in\overline{\varphi}(T_{n})$ with $\gamma\neq\gamma_{h}$. So by (1), $\varphi_{\gamma}=\varphi/(G-T_{n+1},\alpha,\gamma)$ is $T_{n+1}$ mod $\varphi$ and $T_{n+1}$ is an ETT with $n$-rungs under $\varphi_{\gamma}$. Thus $T_{n}$ remains an ETT with (n1)(n-1)-rungs under $\varphi_{\gamma}$. Since $\gamma\neq\gamma_{h}$, we see that there is at most one $(\gamma,\beta)$-path intersecting $T_{n}$ by (A2). Because $P_{1}$, $P_{2}$, and $P_{3}$ are $(\gamma,\beta)$-exit paths of $T_{n+1}$ under $\varphi_{\gamma}$ with exit vertices $b_{1}\prec b_{2}\prec b_{3}$, (2) is established. Depending on whether $\beta$ is $\delta_{h}$ or not, we consider the following two cases.
Case I: $\beta\neq\delta_{h}$.
@@
Note that $P_{3}$ is a $(\gamma,\beta)$-path under $\varphi_{\gamma}$ with $\gamma\in\overline{\varphi}_{\gamma}(b_{3})$. So $f_{i}$ is colored by $\beta$ under both $\varphi$ and $\varphi_{\gamma}$ for $i=1,2,3$. Consider $\sigma=\varphi_{\gamma}/P_{3}$. Note that $\beta\in\overline{\sigma}(b_{3})$, and $f_{i}$ is colored by $\beta$ under $\sigma$ for $i=1,2$. Let $T^{\prime}$ be a closure of $T_{n+1}(b_{3})$ under $\sigma$. Since $\beta\in\overline{\sigma}(b_{3})$, by TAA,
(4) both $f_{1}$ and $f_{2}$ are contained in $G[V(T^{\prime})]$.
We claim that
(5) $\sigma$ is $T_{n}+f_{n}$ mod $\varphi_{\gamma}$, and $T^{\prime}$ is an ETT with (n1)(n-1)-rungs under $\sigma$.
Note that $T^{\prime}$ can be obtained by TAA from $T_{n}+f_{n}$. So by (1) and (3), $T^{\prime}$ is an ETT with (n1)(n-1)-rungs under $\sigma$ once we show that $\sigma$ is $T_{n}+f_{n}$ mod $\varphi_{\gamma}$. Indeed, because $\beta\notin\overline{\varphi}(T_{n+1})$ and $T_{n+1}$ is closed under $\varphi$, we have $\beta\neq\gamma_{h}$. As $\gamma_{h}\in\overline{\varphi}(T_{h})$ and $T_{n+1}$ is elementary under $\varphi$ by (A1), we have $\gamma\neq\gamma_{h}$ by (2). So both $\gamma$ and $\beta$ are different from $\gamma_{h}$ . Recall that we have $\beta\neq\delta_{h}$ in the current case. So if $\gamma\neq\delta_{h}$, then $T_{h}$ has the same $(\gamma_{h},\delta_{h})$-exit vertex under $\sigma$ as $\varphi_{\gamma}$, and thus $\sigma$ is $T_{n}+f_{n}$ mod $\varphi_{\gamma}$ by (A2.5) for $T_{n}$. If $\gamma=\delta_{h}$, then all the $(\gamma_{h},\delta_{h})$-ears of $T_{h}$ under $\varphi$ are contained in $G[V(T_{n+1})]$ by TAA. Thus by the construction of $\varphi_{\gamma}$ and $\sigma$, they are colored the same under $\sigma$ as under $\varphi$. Since $T_{h}$ only has one $(\gamma_{h},\delta_{h})$-exit path under $\varphi$ by (A2), $T_{h}$ still has the same $(\gamma_{h},\delta_{h})$-exit vertex under $\varphi_{\gamma}$ as $\varphi_{\gamma}$, and thus (5) is again established by (A2.5) for $T_{n}$.
@@
Case II: $\beta=\delta_{h}$.
@@
Observe that $P_{j}$ is an $(\eta,\beta)$-path under $\varphi_{\eta}$ with $\eta\in\overline{\varphi}_{\eta}(b_{j})$. Similarly as in the previous case, we let $\sigma=\varphi_{\eta}/P_{j}$ and $T^{\prime}$ be a closure of $T_{n+1}(b_{j})$ under $\sigma$. We claim that
@@
Note that following (7), $T^{\prime}$ is elementary by (A1) and (6), and therefore similarly as earlier, none of $a_{1}$ and $a_{j}$ are in $T^{\prime}$ (see the paragraph before Case II). So under $\sigma$, $P_{1}$ and $P_{j}$ guarantee two $(\beta,\eta)$-paths intersecting $T^{\prime}$, and they provide three $(\beta,\eta)$-exit paths $L_{1},L_{2}$ and $L_{3}$ of $T^{\prime}$, where $L_{1}$, $L_{2}$, and $L_{3}$ produce at least two $(\beta,\eta)$-exit vertices in $T^{\prime}$ not before $b_{j}$ along $\prec$ of $T^{\prime}$ (one of them might be $b_{j}$ itself). Since $\delta_{h}=\beta\in\overline{\sigma}(b_{j})$, $(\sigma,T^{\prime},\beta,\eta,L_{1},L_{2},L_{3})$ provides a counterexample with $Q(T^{\prime})=0$, a contradiction.
@@
We now prove (7) to finish the proof of Lemma 4.1 by assuming otherwise that $T_{h}$ has a different $(\gamma_{h},\delta_{h})$-exit vertex $v_{1}$ under $\sigma$ from the $(\gamma_{h},\delta_{h})$-exit vertex $v_{2}$ under $\varphi_{\eta}$. Note that $\sigma$ is $(T_{n},D_{n},\varphi_{\eta})$-stable as $P_{j}$ is disjoint from $T_{n}$ (see (2)). So $\sigma$ is $T_{h}$ mod $\varphi$ and $T_{h}$ has exactly one $(\gamma_{h},\delta_{h})$-exit vertex $v_{1}$ under $\sigma$ by (A2) for $T_{h}$ with $h\leq n$. We first assume that $v_{1}\prec v_{2}$ along $\prec$ of $T_{n+1}$. In this case, under $\varphi_{\eta}$, $P_{j}$ intersects both the $(\gamma_{h},\delta_{h})$-exit path $P_{ex}$ and a $(\gamma_{h},\delta_{h})$-ear $D$ with a root $v_{1}$ of $T_{h}$. Since $\gamma\neq\gamma_{h}$ and $\eta\neq\gamma_{h}$ (see the sentence before (6)), $P_{ex}$ is still a $(\gamma_{h},\delta_{h})$-exit path and $D$ is still a $(\gamma_{h},\delta_{h})$-ear of $T_{h}$ under $\varphi_{\gamma}$. Let $n^{\prime}$ be the smallest index such that $v_{2}\in T_{n^{\prime}}$ and $T_{n^{\prime}}$ is $\operatorname{RE}$ finished under $\varphi_{\gamma}$. Then $n^{\prime}\leq n$ and $v_{2}\notin T_{h^{\prime}}$, where $h^{\prime}$ is the initial index of $n^{\prime}$. Moreover, under $\varphi_{\gamma}$, $P_{ex}$ is a $(\gamma_{h},\delta_{h})$-exit path of $T_{n^{\prime}}$ by our choice of $n^{\prime}$. Since $T_{n^{\prime}}$ is an ETT with at most (n1)(n-1)-rungs by (1), $v_{1}$ is a root of a $(\gamma_{h},\delta_{h})$-ear $D^{\prime}$ of $T_{n^{\prime}}$ by (A2). Since $v_{1}$ is a root of the $(\gamma_{h},\delta_{h})$-ear $D$ of $T_{h}$ with $h\geq n^{\prime}$, $D$ is a subpath of $D^{\prime}$. So $P$ is a $(\gamma,\delta_{h})$-path intersecting both $P_{ex}$ and $D^{\prime}$. Note that $\gamma\in\overline{\varphi}_{\gamma}(T_{n^{\prime}})$ as $x_{e}$ is the first vertex in $T_{n+1}$ along $\prec$. However, this is a contradiction to the later introduced technical Lemma 4.4 under $\varphi_{\gamma}$ with $T_{n^{\prime}}$ ($n^{\prime}\leq n$) in place of $T_{n+1}$, $D^{\prime}$ in place of $D$, $v_{2}$ in place of $x$, $v_{1}$ in place of $y$, $\gamma$ in place of $\alpha$, and $\gamma_{h},\delta_{h}$ in place of $\gamma,\delta$.
@@
So we may assume $v_{2}\prec v_{1}$. Now in this case, under $\sigma$, we may assume $P_{j}$ intersects both the $(\gamma_{h},\delta_{h})$-exit path $P_{ex}$ of $T_{h}$ with exit $v_{1}$ and a $(\gamma_{h},\delta_{h})$-ear $D$ with root $v_{2}$. Since $T_{n+1}$ is elementary under $\varphi$, $\delta_{h}=\beta\notin\overline{\varphi}(T_{n+1})$, and $\gamma,\eta\neq\gamma_{h}$ (see the sentence before (6)), we observe that:
@@
(8) $\sigma$ is $(T_{h},D_{h},\varphi)$-stable, $T_{n+1}$ is elementary under $\sigma$, $\sigma(P)=\{\gamma,\eta,\delta_{h}\}$, all the $\eta$ colored edges of $P$ are not in $G[V(T_{n+1})]$, all the $\gamma$ colored edges of $P$ are in $G[V(T_{n+1})]$, $\delta_{h}\in\overline{\sigma}(b_{j})$, two ends of $P$ miss either $\delta_{h}$ or $\eta$, and exactly one boundary edge of $T_{n+1}$ is colored by $\eta$ which is $f_{j}$ under $\sigma$.